OpenStudy (anonymous):

The radius of a circle of area A and circumference C is doubled. a) find the new area of the circle in terms of A b) find the new circumference of the circle in terms of C Please explain.

OpenStudy (anonymous):

A1 = pi (r1)^2 A2 = pi (2r1)^2 c1 = 2 pi (r1) c2 = 2 pi (2 r1) the rest is yours

OpenStudy (anonymous):

I got the same answer but they told me that the answer does not depend on pi and r

OpenStudy (anonymous):

You make the ratios A2/A1 and c2/c1 to answer the question as given. you end up with (2 r1)^2/(r1)^2 and (2r1/r1) to simplify further.

OpenStudy (anonymous):

thanks for the help!

OpenStudy (anonymous):

You are quite welcome.